While standing on the ground, you want to throw a textbook straight up to a classmate standing by a window 12.p m above you. With what velocity must you throw the book so that the maximum height it reaches is the height of the window?
1
Expert's answer
2019-10-04T12:55:28-0400

Solution:

Let's use the law of conservation of energy. Potential energy of textbook on the ground is equal to zero, and kinetic energy of the textbook when it reaches the maximum height is also equal to zero. So kinetic energy of the book on the ground is equal to it's potential energy in the highest point:


"\frac{m\upsilon^2}{2}=mgh"


"\upsilon=\sqrt{2gh}=\sqrt{2\cdot{}9.8\cdot{}12}=15" m/s.


Answer:

15 m\s.